# DRWKV: Focusing on Object Edges for Low-Light Image Enhancement

### v1

@@ Abstract
Low-light image enhancement remains a challenging task, particularly in preserving object edge continuity and fine structural details under extreme illumination degradation. In this paper, we propose a novel model, DRWKV (Detailed Receptance Weighted Key Value), which integrates our proposed Global Edge Retinex (GER) theory, enabling effective decoupling of illumination and edge structures for enhanced edge fidelity. Secondly, we introduce Evolving WKV Attention, a spiral-scanning mechanism that captures spatial edge continuity and models irregular structures more effectively. Thirdly, we design the Bilateral Spectrum Aligner (Bi-SAB) and a tailored MS²-Loss to jointly align luminance and chrominance features, improving visual naturalness and mitigating artifacts. Extensive experiments on five LLIE benchmarks demonstrate that DRWKV achieves leading performance in PSNR, SSIM, and NIQE while maintaining low computational complexity. Furthermore, DRWKV enhances downstream performance in low-light multi-object tracking tasks, validating its generalization capabilities.
@@ III-B Deep Detail Mining
Here, $R$ strips off spatially heterogeneous noise via $I-N$ and neutralizes illumination interference through $L+\epsilon$, providing a ”clean” benchmark for detail enhancement.
@@ III-C Bilateral Spectrum Aligner
2:  $F_{Q}:{\color[rgb]{0,0,1}(B,C_{i},H_{i},W_{i})}\leftarrow\text{Conv}_{3\times 3}(F_{1})$
3:  $F_{K}:{\color[rgb]{0,0,1}(B,C_{i},H_{i},W_{i})}\leftarrow\text{Conv}_{3\times 3}(F_{1}^{{}^{\prime}})$
4:  $F_{V}:{\color[rgb]{0,0,1}(B,C_{1},H_{i},W_{i})}\leftarrow\text{Conv}_{3\times 3}(F_{1}^{{}^{\prime}})$
@@
8:  $Att_{out}:{\color[rgb]{0,0,1}(B,C_{1},H_{i},W_{i})}\leftarrow F_{attn}\times F_{K}$
9:  $F_{DA}:{\color[rgb]{0,0,1}(B,C_{1},H_{i},W_{i})}\leftarrow\lambda\cdot Att_{out}+F_{V},\lambda=0.2$
10:  $DA_{out}:{\color[rgb]{0,0,1}(B,C_{1},H_{i},W_{i})}\leftarrow F_{attn}\times F_{Q}+DA_{out}$
@@ III-D Loss Function
[…]
@@ IV-C Ablation Studies
[…]
@@ VII-B1 Evolving Spatial Mix
In this context, (){$(*)\in\{\textit{R},\textit{K},\textit{V}\}$ denotes three types of interpolation operations performed on $X$ and $X^{\dagger}$ These operations are governed by $\mu$ and are utilized for the subsequent computation of R, K, and V. Here, h, w, and C represent the height, width, and number of channels, respectively.
@@
where $\textit{where }wkv=\textit{EV-WKV}(K_{s},V_{s}),$ (25)
@@ VII-B2 Channel Mix
[…]
@@
where $\textit{where }V_{c}=SR((O_{s}+(1-\mu_{(k)})O_{s}^{\dagger})W_{K})W_{V}$ (28)
@@ VII-C3 Scharr Edge Enhancement (SEE)
Metric RetinexNet Zero-DCE RRDNet RetinexFormer RetinexMamba MambaIR DRWKV
@@
To address the above issues, we propose a Scharr Edge Enhancement (SEE) module. By integrating the prior knowledge of traditional image processing with the feature learning capability of deep learning, it provides an effective solution for structure-aware enhancement in low-light scenarios. Specifically, given the input feature $F\in\mathbb{R}^{C\times H\times W}$, we first perform channel-wise filtering operations using pre-fixed Scharr convolution kernels and convolution weights $W_{x}$,$W_{y}$ to simulate the gradient extraction process in traditional image processing, obtaining the x-direction gradient feature $E_{x}\in\mathbb{R}^{C\times H\times W}$ and y-direction gradient feature $E_{y}\in\mathbb{R}^{C\times H\times W}$ respectively.

### v2

@@ Abstract
Low-light image enhancement remains a challenging task, particularly in preserving object edge continuity and fine structural details under extreme illumination degradation. In this paper, we propose a novel model, DRWKV (Detailed Receptance Weighted Key Value), which integrates our proposed Global Edge Retinex (GER) theory, enabling effective decoupling of illumination and edge structures for enhanced edge fidelity. Secondly, we introduce Evolving WKV Attention, a spiral-scanning mechanism that captures spatial edge continuity and models irregular structures more effectively. Thirdly, we design the Bilateral Spectrum Aligner (Bi-SAB) and a tailored MS²-Loss to jointly align luminance and chrominance features, improving visual naturalness and mitigating artifacts. Extensive experiments on five LLIE benchmarks demonstrate that DRWKV achieves leading performance in PSNR, SSIM, and NIQE while maintaining low computational complexity. Furthermore, DRWKV enhances downstream performance in low-light multi-object tracking tasks, validating its generalization capabilities.The code are available at: https://github.com/JackBaixue/DRWKV
@@ III-B Deep Detail Mining
Here, $R$ strips off spatially heterogeneous noise via INI-N and neutralizes illumination interference through $L+\epsilon$, providing a ”clean” benchmark for detail enhancement.
@@ III-C Bilateral Spectrum Aligner
[…]
@@
[…]
@@ III-D Loss Function
The newly defined $MS^{2}\text{-Loss}$ focuses on five key attributes of the model: structure, edge, illumination, artifact, and weight.
@@ IV-C Ablation Studies
Validation (5): Replacing the SSIM+L1 loss with the proposed $MS^{2}\text{-Loss}$ specifically optimizes the training process, with SSIM and PSNR optimized to 0.832 and 24.12 dB, respectively.
@@ VII-B1 Evolving Spatial Mix
In this context, $(*)\in\{\textit{R},\textit{K},\textit{V}\}$ denotes three types of interpolation operations performed on $X$ and $X^{\dagger}$ . These operations are governed by $\mu$ and are utilized for the subsequent computation of R, K, and V. Here, h, w, and C represent the height, width, and number of channels, respectively.
@@
$\textit{where }wkv=\textit{EV-WKV}(K_{s},V_{s}),$ (25)
@@ VII-B2 Channel Mix
Here, $\textit{Q-Shift}_{R}$ and $\textit{Q-Shift}_{K}$ dynamically interpolate between the original feature $O_{s}$ and its shifted version $O^{\dagger}_{s}$ using learnable parameters $\mu_{(R)}$ and $\mu_{(K)}$. This design not only preserves the spatial structural information in the original features but also introduces cross-position channel dependencies, enabling the model to adaptively focus on long-range correlations among different channels.
@@
$\textit{where }V_{c}=SR((O_{s}+(1-\mu_{(k)})O_{s}^{\dagger})W_{K})W_{V}$ (28)
@@ VII-C3 Scharr Edge Enhancement (SEE)
Metric
RetinexNet Zero-DCE RRDNet RetinexFormer RetinexMamba MambaIR DRWKV
@@
To address the above issues, we propose a Scharr Edge Enhancement (SEE) module. By integrating the prior knowledge of traditional image processing with the feature learning capability of deep learning, it provides an effective solution for structure-aware enhancement in low-light scenarios. Specifically, given the input feature $F\in\mathbb{R}^{C\times H\times W}$, we first perform channel-wise filtering operations using pre-fixed Scharr convolution kernels and convolution weights $W_{x}$$W_{y}$ to simulate the gradient extraction process in traditional image processing, obtaining the x-direction gradient feature $E_{x}\in\mathbb{R}^{C\times H\times W}$ and y-direction gradient feature $E_{y}\in\mathbb{R}^{C\times H\times W}$ respectively.